# A shared-revenue Bertrand game

### v1

@@ 1.2. Overview
In this work we provide a game theoretic model for reasoning about the interaction between agents in scenarios such as these with a variation of the Bertrand pricing game (Bertrand, 1883). For concreteness, we will focus our intuition and analysis on scenario (A) from Section 1.1, though our model can explain phenomena in domains far beyond retail sales. In accordance with scenario (A), the “shared-revenue Bertrand game” is played between a retailer, and an independent goods/services provider which we will refer to as the independent seller (seller). Both players sell an identical good/service, with each agent having a different cost. Under the revenue sharing program, whenever the independent seller fulfills demand, a portion of the revenue, called the referral fee, is transferred from the independent seller to the retailer. In Section 2, we set up the game in more detail, then in Section 3 we identify the Nash equilibria of this game. In Section 4, we interpret and analyze the equilibria of the shared-revenue Bertrand game in depth with different costs and fee structures.
The referral fee is a constant parameter of the shared-revenue Bertrand game of Section 2. However, to make the game more realistic, in Section 5 we introduce the “fee optimization game,” which models the referral fee as an action for the retailer that is chosen before playing the shared-revenue Bertrand game. Finally, in Section 6, we introduce the “outside option game”: after observing the referral fee chosen by the retailer, the independent seller can choose to sell on their own instead of through the retailer.
@@
The equilibria of the “fee optimization game” reveal that the retailer always prefers to allow the independent seller to fulfill demand in equilibrium, setting the referral fee strategically so as to maximize their payoff. There is a sweet spot for the optimal fee such that it is low enough to stimulate economic activity while high enough to maximize payoff. In the fee optimization game of Section 5, we still observe some unrealistic equilibria in which the retailer sets such a high fee that the independent seller must fulfill demand at their effective cost (thus achieving no payoff), which leads us to consider the “outside option game” in Section 6. If the independent seller’s threat of selling on their own is credible, the unrealistic equilibria of Section 5 will no longer be observed, as the retailer must set the referral fee lower to ensure the independent seller doesn’t exercise their outside option. Importantly, our analysis demonstrates that the retailer will frequently be incentivized to set a fee under which both the independent seller increases its payoff compared to selling independently, and the price to the customer is decreased relative to the price either the retailer or the independent seller would have set independently. Our results therefore suggest that such a revenue sharing mechanism can not only increase social welfare, but also simultaneously improve outcomes for both players as well as consumers.
In each of the models, we examine if the revenue sharing program increases the retailer’s payoff compared to simply fulfilling demand on their own, and how it affects the end cost to the customer.
@@ 1.3. Related work
Our model is an extension of the standard Bertrand game (Bertrand, 1883) in which two sellers of an identical product with the same cost set their prices independently. As we justify in Appendix B, we consider the regime in which the sellers’ costs are asymmetric, which is a natural extension of the standard Bertrand game, studied for instance in (Blume, 2003; Kartik, 2011; Demuynck, 2019).
One field in which revenue sharing programs as described in Section 1.1 have been studied is in the independent contractor literature. Much of this literature analyzes relationships with independent contractors through the framework of principal agent theory (Coats, 2002; Chang, 2014). The revenue-sharing program can also be seen as a form of “make-or-buy” decision for the retailer (Coase, 1937; Williamson, 1975; Tadelis, 2002). Slightly less obviously, another field in which revenue sharing programs have been studied is optimal taxation theory, where the referral fee can be viewed as the retailer imposing a tax on the independent seller. Viewed in this light, our results about optimal referral fees resonate strongly with optimal taxation theories such as (Laffer, 2004; Stern, 1976; Stiglitz, 1987).
@@ 2. Preliminaries
Our model is a game in which two players – the retailer ($r$), and the independent seller ($s$), sell an identical product. The actions in the game are the price each player chooses to charge, and we assume the players move simultaneously. Formally, we define the following game:
Shared-Revenue Bertrand Game
@@
Retailer’s cost $c_{r}$, seller’s cost $c_{s}$, referral fee $\alpha$
@@
(1) $\bar{\pi}_{r}$ $=(\bar{p}_{r}-\bar{c}_{r})\bar{q}_{r}+\alpha\bar{p}_{s}\bar{q}_{s}$
(2) $\bar{\pi}_{s}$ $=(\bar{p}_{s}-\bar{c}_{s}-\alpha\bar{p}_{s})\bar{q}_{s}$
where $\bar{p}$, $\bar{c}$, $\bar{q}$ and $\alpha$ denote prices, costs, quantities sold and the referral fee, respectively, and subscripts denote the players. This game can be viewed as a modification of the Bertrand game, where the players have different costs, and a portion of the independent seller’s revenue is shared with the retailer as part of the service provided by the retailer.
To allow for a concrete mathematical analysis we assume demand is linear in price, and that the agent with the lower price fulfills all demand. When both players set the same price, we assume they split the market where the retailer fulfills $\beta\in[0,1]$ of the demand and the independent seller fulfills the rest. This assumption is made for mathematical completeness, although in Proposition 1 we show that no split market equilibrium exists, making the choice of $\beta$ irrelevant. Concretely, the retailer’s demand curve is given by
(3) $\bar{q}_{r}(\bar{p}_{r},\bar{p}_{s})=\begin{cases}Q_{0}-K^{(q)}\bar{p}_{r}&\bar{p}_{r}<\bar{p}_{s}\\ \beta(Q_{0}-K^{(q)}\bar{p}_{r})&\bar{p}_{r}=\bar{p}_{s}\\ 0&\bar{p}_{r}>\bar{p}_{s}\end{cases}$
where $Q_{0}$ and $K^{(q)}$ are constant parameters. The demand curve for the independent seller is analogous, but $\beta$ is replaced by $(1-\beta)$.
To reduce the number of parameters in the model we will use the following normalization, and throughout the rest of the paper present the results in terms of unitless variables
(4) $p=\bar{p}\frac{K^{(q)}}{Q_{0}};\quad c=\bar{c}\frac{K^{(q)}}{Q_{0}};\quad q=\bar{q}\frac{1}{Q_{0}};\quad\pi=\bar{\pi}\frac{K^{(q)}}{Q_{0}^{2}}.$
Under this normalization, the demand at price $p=0$ is 1, and the demand at price $p=1$ is 00. Equations (1) and (2) remain unchanged if we remove bars, and (3) reduces to
(5) $q_{r}(p_{r},p_{s})=\begin{cases}1-p_{r}&p_{r}<p_{s}\\ \beta(1-p_{r})&p_{r}=p_{s}\\ 0&p_{r}>p_{s}\end{cases}$
@@ 2.1. Necessary assumptions
    There exists a price at which both the retailer and independent seller could achieve a nonzero payoff when fulfilling demand ($10<c_{s},c_{r}<$ 1).
@@
2.2. Important prices and payoffs
We now introduce several important prices and corresponding payoffs which will be used throughout this paper. We also include a full notation table in Appendix A, and include derivations of the key prices and fees in Appendix D.
@@ 2.2. Important prices and payoffs
(6) $\pi_{r,r}(p_{r})=(p_{r}-c_{r})(1-p_{r})\ ;$ $\pi_{s,r}(p_{r})=0$
(7) $\pi_{r,s}(p_{s})=\alpha p_{s}(1-p_{s})\ ;$ $\pi_{s,s}(p_{s})=((1-\alpha)p_{s}-c_{s})(1-p_{s})$
In Figure 1 we plot the payoffs of the players as a function of price for two choices of game parameters $(c_{r},c_{s},\alpha)$. Solid and dashed lines denote the players’ payoff when the retailer or independent seller fulfill all demand, respectively, while blue and red lines denote the payoffs to the retailer and the independent seller, respectively.
Figure 1. Typical payoff curves and important price points
Figure 1 also illustrates five important price points and their corresponding payoffs which will be used throughout this paper. We denote by $p_{r^{*}}$ and $p_{s^{*}}$ the prices each player would set to maximize their own payoff if they fulfilled all demand, and call these the retailer’s or independent seller’s optimal selling price. Differentiating the players’ single agent payoffs $\pi_{r,r}$ and $\pi_{s,s}$ and equating to zero in (36) and (40) gives the values of these prices and their corresponding payoffs –
(8) $p_{r^{*}}\equiv\frac{1+c_{r}}{2};$ $\pi_{r,r}(p_{r^{*}})=\frac{(1-c_{r})^{2}}{4}$
(9) $p_{s^{*}}\equiv\frac{1+\frac{c_{s}}{1-\alpha}}{2};$ $\pi_{s,s}(p_{s^{*}})=\begin{cases}\frac{1-\alpha}{4}\left(1-\frac{c_{s}}{1-\alpha}\right)^{2}&\alpha<1-c_{s}\\ 0&\alpha\geq 1-c_{s}\end{cases}$
@@
(10) $p_{r_{\text{ind}}}\equiv\frac{c_{r}}{1-\alpha};$ $\pi_{r,r}(p_{r_{\text{ind}}})=\begin{cases}\frac{\alpha c_{r}}{1-\alpha}\left(1-\frac{c_{r}}{1-\alpha}\right)&\alpha<1-c_{r}\\ 0&\alpha\geq 1-c_{r}\end{cases}$
(11) $p_{s_{\text{ind}}}\equiv\frac{c_{s}}{1-\alpha};$ $\pi_{s,s}(p_{s_{\text{ind}}})=0$
We first note that $\pi_{s,s}(p_{s_{\text{ind}}})=0$, and therefore the independent seller’s indifference price is also its breakeven price.
More interestingly, while the retailer’s breakeven price is still its cost, $c_{r}$, its indifference price is scaled by $(1-\alpha)^{-1}$, but at that price its payoff is nonzero. This is the main feature which differentiates the shared-revenue Bertrand game from the classical Bertrand game, as the retailer can choose to fulfill demand on its own or allow the independent seller to fulfill demand, making revenue from the fee. As a result, this game gives rise to a more complex equilibrium structure which depends on the fees and costs of the players. Note also that because we assume $c_{s}<c_{r}$ (Section 2.1), we have $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$.
Next, we introduce the price at which the retailer’s payoff when the independent seller fulfills demand is equal to its optimal payoff had it fulfilled demand on its own, i.e.
$\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$
(12) $p^{\dagger}\equiv\frac{1+\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}$
$p^{\dagger}$ is the solution of a quadratic equation which we derive in (58), and we define it as the larger of the two possible solutions. Furthermore, $p^{\dagger}$ is only well-defined when $\alpha\geq(1-c_{r})^{2}$, and therefore we only show it in the right subplot of Figure 1.
@@ 3.1. Market split
Let $(p_{r},p_{s})=(p,p)$ be an equilibrium price configuration, with the equilibrium quantities sold by each player $q_{r}=\beta(1-p)$ and $q_{s}=(1-\beta)(1-p)$. Then $\beta\in\{0,1\}$.
@@
To prove the Proposition we show that when the market is split, the retailer’s payoff is a weighted average of $\pi_{r,r}$ and $\pi_{r,s}$, and one of them is larger than the split market payoff. Thus, the retailer can always increase their payoff by either lowering or raising the price by an infinitesimal amount, fulfilling demand themselves or allowing the independent seller to do so respectively. See Appendix C for the complete proof. ∎
@@ 3.2. Nash equilibrium derivation
We acknowledge that it’s still not clear what the equilibrium interval of prices look like for each possible choice of $\alpha$, namely what lower and upper bounds are attained for each possible referral fee – this warrants a longer discussion which we defer to Section 3.3. For instance, $p^{\dagger}$ is not well-defined for small $\alpha$, but we will soon see in (20) that whenever $\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$ is nonempty, $\alpha$ is high enough for $p^{\dagger}$ to be well-defined. For now, simply keep in mind that (13) is sufficient only for shared-price Nash equilibria where the independent seller fulfills demand.
@@ 3.3. Important fees for transitions between Nash equilibria
Consider first the left interval of (13), $p\in\left[p_{s_{\text{ind}}},\min\{p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$. We know that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always. We have $p_{s_{\text{ind}}}\leq p_{s^{*}}=\frac{1}{2}(1+p_{s_{\text{ind}}})$ so long as $p_{s_{\text{ind}}}\leq 1$, which occurs when $\alpha\leq 1-c_{s}$. The condition for $p_{s_{\text{ind}}}\leq p_{r^{*}}$ is not quite as clear, so we introduce the retailer optimal price feasibility fee $\alpha_{r^{*}}$ (derived in (46)) –
(15) $p_{s_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{r^{*}}\ ;$ where $\text{where }\alpha_{r^{*}}\equiv 1-\frac{2c_{s}}{1+c_{r}}$
The name comes from the fact that $p_{r^{*}}\leq p_{s_{\text{ind}}}$ is necessary for any equilibrium where the retailer fulfills demand at their optimal price, which we prove in Appendix D.3. Note also that $\alpha_{r^{*}}\leq 1-c_{s}$, which we show in (52). Thus, the left interval of (13) is nonempty if and only if $\alpha\leq\alpha_{r^{*}}$.
@@
(16) $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}\ ;$ where $\text{where }\alpha_{s^{*},r^{*}}\equiv\frac{c_{r}-c_{s}}{c_{r}}$
(17) $p_{r_{\text{ind}}}\leq p_{s^{*}}\iff\alpha\leq\alpha_{s^{*}}\ ;$ where $\text{where }\alpha_{s^{*}}\equiv 1-2c_{r}+c_{s}$
(18) $p_{r_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}\ ;$ where $\text{where }\alpha^{\dagger}_{r}\equiv\frac{1-c_{r}}{1+c_{r}}$
@@
(19) $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}\iff\alpha^{\dagger}_{r}\leq\alpha_{s^{*},r^{*}}\iff c_{s}\leq c_{s^{*}}\ ;$ where $\text{where }c_{s^{*}}\equiv\frac{2c_{r}^{2}}{1+c_{r}}$
Finally, let’s turn our attention to the right interval of (13), $p\in\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$. Fortunately, we have already determined the relative ordering of each of the prices in the right interval, except for $p^{\dagger}$. For these conditions involving $p^{\dagger}$, we introduce the $p^{\dagger}$ relevance fees for the retailer and seller (derived in (59) and (62)):
(20) $p^{\dagger}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}\ ;$ where $\text{where }\alpha^{\dagger}_{r}\equiv\frac{1-c_{r}}{1+c_{r}}$
@@
Note that we already introduced $\alpha^{\dagger}_{r}$ in (18). The notation arises because we will show that $\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}$ is necessary for any equilibrium where demand is fulfilled at $p^{\dagger}$. We don’t include a closed form here for $\alpha^{\dagger}_{s}$ because it is uninsightful, but we remark that $\alpha^{\dagger}_{s}\in[\alpha_{r^{*}},1-c_{s}]$ and that a closed form is in (62). Finally, the relative ordering of $p_{s^{*}}$ and $p^{\dagger}$ is complex, and there unfortunately is no special $\hat{\alpha}$ for which $p_{s^{*}}\leq p^{\dagger}\iff\alpha\leq\hat{\alpha}$. Thus, we will leave that condition unsimplified for now.

### v2

@@ 1.2. Overview
In this work we provide a game theoretic model for reasoning about the interaction between agents in scenarios such as these with a variation of the Bertrand pricing game (Bertrand, 1883). For concreteness, we will focus our intuition and analysis on scenario (A) from Section 1.1, though our model can explain phenomena in domains far beyond retail sales. In accordance with scenario (A), the “shared-revenue Bertrand game” is played between a retailer (r), and an independent goods/services provider which we will refer to as the independent seller (seller, s). Both players sell an identical good/service, with each agent having a different cost. Under the revenue sharing program, whenever the independent seller fulfills demand, a portion of the revenue, called the referral fee $(\alpha)$, is transferred from the seller to the retailer. Specifically, the game is as follows:
Shared-Revenue Bertrand Game
Parameters
Demand curve $q(p)$, retailer’s cost $c_{r}$, seller’s cost $c_{s}$ [^1] The leaving subgame of Section 6.1 is additionally parameterized by the outside option cost differential $\delta\in(0,c_{r}-c_{s})$.
Gameplay
  1. (1)
    The retailer chooses a referral fee $\alpha\in(0,1)$ (Section 5)
  2. (2)
    The independent seller chooses one of two options (Section 6):
    Stay in the retailer’s revenue sharing program
    Leave the retailer’s revenue sharing program and sell on their own
  3. (3)
    The retailer and seller choose their selling prices:
    If the independent seller chose to stay, they play a staying subgame (Sections 2-4)
    If the independent seller chose to leave, they play a leaving subgame (Section 6.1)
The primary contribution of this work is in the half of the game tree where the independent seller stays in the revenue sharing program and the firms compete with coupled payoff functions. In Section 2, we set up the staying game in more detail, then in Section 3 we solve for the Nash equilibria of this subgame. In Section 4, we interpret and analyze the equilibria of the staying subgame in depth, examining the effects of the revenue sharing program on the firms’ payoffs and the end price to the customer.
@@
The referral fee is a constant parameter of the staying subgame of Section 2. In Section 5 we proceed by analyzing the “fee optimization game,” which is the same as the the shared-revenue Bertrand game except the independent seller does not have the option to leave the revenue sharing program. The equilibria of the fee optimization game reveal that the retailer always prefers to allow the independent seller to fulfill demand in equilibrium, setting the referral fee strategically so as to maximize their payoff. There is a sweet spot for the optimal fee such that it is low enough to stimulate economic activity while high enough to maximize payoff. However, we still observe some unrealistic equilibria in which the retailer sets such a high fee that the independent seller must fulfill demand at their effective cost.
Finally, in Section 6, we solve the entire shared-revenue Bertrand game, reintroducing the ability for the independent seller to leave the revenue sharing program after observing the retailer’s choice of referral fee. In Section 6.1, we choose the leaving subgame to be one of standard Bertrand competition with asymmetric costs. Since the literature on direct price competition is extensive, we do not focus closely on choosing the leaving subgame (Bertrand, 1883; Shubik and Levitan, 1980; Tirole, 1988). With this, if the independent seller’s threat of selling on their own is credible, the unrealistic equilibria of the fee optimization game will no longer be observed. The retailer must instead set the referral fee low enough to ensure the independent seller doesn’t exercise their outside option.
Importantly, our analysis demonstrates that the retailer will often be incentivized to set a fee at which both the independent seller increases its payoff compared to selling independently, and the price to the customer is decreased relative to the price either the retailer or the independent seller would have set independently. Our results therefore suggest that such a revenue sharing mechanism can not only increase social welfare, but also simultaneously improve outcomes for both players as well as consumers.
@@ 1.3. Related work
Our model is an extension of the standard Bertrand game (Bertrand, 1883) in which two sellers of an identical product with the same cost set their prices independently. As we justify in Appendix B, we consider the regime in which the sellers’ costs are asymmetric, which is a natural extension of the standard Bertrand game, studied for instance in (Blume, 2003; Kartik, 2011; Demuynck, 2019). Prior works have pursued other modifications to the Bertrand game to better describe platform competition, for instance in (Rochet and Tirole, 2003; Angelini et al., 2024; Armstrong, 2006).
One field in which revenue sharing programs as described in Section 1.1 have been studied is in the independent contractor literature. Much of this literature analyzes relationships with independent contractors through the framework of principal agent theory (Coats, 2002; Chang, 2014). The revenue-sharing program can also be seen as a form of “make-or-buy” decision for the retailer (Coase, 1937; Williamson, 1975; Tadelis, 2002). Some revenue sharing programs are implemented via two-part tariffs which involve a fixed base fee alongside the per-unit fee (Yin, 2004; Griva and Vettas, 2015). A final field in which revenue sharing programs have been studied is optimal taxation theory, where the referral fee can be viewed as the retailer imposing a tax on the independent seller. Viewed in this light, our results about optimal referral fees resonate strongly with optimal taxation theories such as (Laffer, 2004; Stern, 1976; Stiglitz, 1987).
@@ 2. Preliminaries
The core staying subgame is one in which two players – the retailer ($r$), and the independent seller ($s$), sell an identical product. The actions in the game are the price each player chooses to charge, and we assume the players move simultaneously. Formally, we define the following game:
Staying Subgame
@@
Demand curve $q(p)$, referral fee $\alpha$, retailer’s cost $c_{r}$, seller’s cost $c_{s}$
@@
(1) $\overline{\pi}_{r}$ $=(\overline{p}_{r}-\overline{c}_{r})\overline{q}_{r}+\alpha\overline{p}_{s}\overline{q}_{s}$
(2) $\overline{\pi}_{s}$ $=(\overline{p}_{s}-\overline{c}_{s}-\alpha\overline{p}_{s})\overline{q}_{s}$
where $\overline{p}$, $\overline{c}$, $\overline{q}$ and $\alpha$ denote prices, costs, quantities sold and the referral fee, respectively, and subscripts denote the players. This game can be viewed as a modification of the Bertrand game, where the players have different costs, and a portion of the independent seller’s revenue is shared with the retailer as part of the service provided by the retailer.
We will suppose that the demand curve $\overline{q}(\overline{p})$ is strictly concave, compactly supported, monotonically decreasing, twice-differentiable, and nonnegative [^2] The only “real” assumptions here are concavity and continuity. Compact support is imposed for convenience so we can normalize to $[0,1]$. If monotonicity and/or strictness are violated, the key prices and fees may become nonunique (connected intervals). If $q$ is only once-differentiable, we can replace derivatives with subderivatives.. When both players set the same price, we also assume they split the market where the retailer fulfills $\beta\in[0,1]$ of the demand and the independent seller fulfills the rest. This assumption is made for mathematical completeness, although in Proposition 1 we show that no split market equilibrium exists, making the choice of $\beta$ irrelevant. Concretely, the retailer’s demand curve is given by
(3) $\overline{q}_{r}(\overline{p}_{r},\overline{p}_{s})=\begin{cases}\overline{q}(\overline{p}_{r})&\overline{p}_{r}<\overline{p}_{s}\\ \beta\overline{q}(\overline{p}_{r})&\overline{p}_{r}=\overline{p}_{s}\\ 0&\overline{p}_{r}>\overline{p}_{s}\end{cases}$
The demand curve for the independent seller is analogous, but $\beta$ is replaced by $(1-\beta)$. Because $q$ is compactly supported, there exists a maximum price $\overline{p}_{\max}\equiv\inf\{\overline{p}\mid\overline{q}(\overline{p})=0\}$. Throughout the rest of the paper we present the results in terms of unitless variables
(4) $p=\frac{\overline{p}}{\overline{p}_{\max}}\ ;\quad c=\frac{\overline{c}}{\overline{p}_{\max}}\ ;\quad q(p)=\frac{\overline{q}(p)}{\overline{q}(0)}\ ;\quad\pi(p_{r},p_{s})=\frac{\overline{\pi}(p_{r},p_{s})}{\overline{q}(0)\overline{p}_{\max}}\ .$
Under this normalization, the demand at price $p=0$ is 1, and the demand at price $p=1$ is 0. Equations (1), (2), and (3) remain unchanged if we remove bars. Note that under this normalization, $q$ inherits the properties of $\overline{q}$ (concavity, monotonicity, differentiability, and nonnegativity).
@@ 2.1. Necessary assumptions
    There exists a price at which both the retailer and independent seller could achieve a nonzero payoff when fulfilling demand ($0<c_{s},c_{r}<1$).
@@
2.2. Important payoffs and prices
We now introduce several important prices which will be used throughout this paper. We also include a full notation table in Appendix A, and include derivations of the key prices and fees in Appendix C.
@@ 2.2. Important payoffs and prices
(5) $\pi_{r,r}(p_{r})=(p_{r}-c_{r})q(p_{r})\ ;$ $\pi_{s,r}(p_{r})=0$
(6) $\pi_{r,s}(p_{s})=\alpha p_{s}q(p_{s})\ ;$ $\pi_{s,s}(p_{s})=((1-\alpha)p_{s}-c_{s})q(p_{s})$
In Figure 1 we plot the payoffs of the players as a function of price for two choices of game parameters $(c_{r},c_{s},\alpha)$ in the special case that q(p)=1pq(p)=1-p is linear. All subsequent figures throughout the paper will also use this linear demand curve for consistency. Solid and dashed lines denote the players’ payoffs when the retailer or independent seller fulfill all demand, respectively, while blue and red lines denote the payoffs to the retailer and the independent seller, respectively.
Figure 1. Sample payoff curves and important price points for linear demand curve q(p)=1pq(p)=1-p
Figure 1 also illustrates five important price points and their corresponding payoffs which will be used throughout this paper. We denote by $p_{r^{*}}$ and $p_{s^{*}}$ the prices each player would set to maximize their own payoff if they fulfilled all demand, and call these the retailer’s or independent seller’s optimal selling price. Differentiating the players’ single agent payoffs $\pi_{r,r}$ and $\pi_{s,s}$ and equating to zero gives first-order conditions for these prices in (32) and (33).
@@
(7) $p_{r_{\text{ind}}}\equiv\frac{c_{r}}{1-\alpha}\ ;$ $p_{s_{\text{ind}}}\equiv\frac{c_{s}}{1-\alpha}$
We first note that $\pi_{s,s}(p_{s_{\text{ind}}})=0$, and therefore the independent seller’s indifference price is also its breakeven price. More interestingly, while the retailer’s breakeven price is still its cost, $c_{r}$, its indifference price is scaled by $(1-\alpha)^{-1}$, but at that price its payoff is nonzero. This is the main feature which differentiates the shared-revenue Bertrand game from the classical Bertrand game, as the retailer can choose to fulfill demand on its own or allow the independent seller to fulfill demand, making revenue from the fee. As a result, this game gives rise to a more complex equilibrium structure which depends on the fees and costs of the players. Note also that because we assume $c_{s}<c_{r}$ (Section 2.1), we have $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$.
Next, we introduce the price at which the retailer’s payoff when the independent seller fulfills demand is equal to its optimal payoff had it fulfilled demand on its own, i.e. $p^{\dagger}$ is the largest price such that
(8) $\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$
Note that $p^{\dagger}$ is only well-defined when $\alpha$ is sufficiently large; for instance it is not well-defined for the parameters in the left subplot of Figure 1 by inspection.
@@ 3.1. Market split
Let $(p_{r},p_{s})=(p,p)$ be an equilibrium price configuration, with the equilibrium quantities sold by each player $q_{r}=\beta q(p)$ and $q_{s}=(1-\beta)q(p)$. Then $\beta\in\{0,1\}$.
@@
To prove the Proposition we show that when the market is split, the retailer’s payoff is a weighted average of $\pi_{r,r}$ and $\pi_{r,s}$, and one of them is larger than the split market payoff. Thus, the retailer can always increase their payoff by either lowering or raising the price by an infinitesimal amount, fulfilling demand themselves or allowing the independent seller to do so respectively. See Appendix C.1 for the complete proof. ∎
@@ 3.2. Nash equilibrium derivation
We acknowledge that it’s still not clear what the equilibrium interval of prices look like for each possible choice of $\alpha$, namely what lower and upper bounds are attained for each possible referral fee – this warrants a longer discussion which we defer to Section 3.3. For instance, $p^{\dagger}$ is not well-defined for small $\alpha$, but we will see in (16) and Lemma 4 that whenever $\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$ is nonempty, $\alpha$ is high enough for $p^{\dagger}$ to be well-defined. For now, simply keep in mind that (9) is sufficient only for shared-price Nash equilibria where the independent seller fulfills demand.
@@ 3.3. Important fees for transitions between Nash equilibria
Consider first the left interval of (9), $p\in\left[p_{s_{\text{ind}}},\min\{p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$. We know that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always. We have $p_{s_{\text{ind}}}\leq p_{s^{*}}$ so long as $p_{s_{\text{ind}}}\leq 1$, which occurs when $\alpha\leq 1-c_{s}$. The condition for $p_{s_{\text{ind}}}\leq p_{r^{*}}$ is not quite as clear, so we introduce the retailer optimal price feasibility fee $\alpha_{r^{*}}$ (derived in (36)) –
(11) $p_{s_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{r^{*}}\ ;$ $\text{where }\alpha_{r^{*}}\equiv 1-\frac{c_{s}}{p_{r^{*}}}$
The name comes from the fact that $p_{r^{*}}\leq p_{s_{\text{ind}}}$ is necessary for any equilibrium where the retailer fulfills demand at their optimal price, which we prove in Appendix D.3. Note also that $\alpha_{r^{*}}\leq 1-c_{s}$ since $p_{r^{*}}\in(0,1)$. Thus, the left interval of (9) is nonempty if and only if $\alpha\leq\alpha_{r^{*}}$.
@@
(12) $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}\ ;$ $\text{where }\alpha_{s^{*},r^{*}}\equiv 1-\frac{c_{s}}{c_{r}}$
(13) $p_{r_{\text{ind}}}\leq p_{s^{*}}\iff\alpha\leq\alpha_{s^{*}}\ ;$ $\text{where }\alpha_{s^{*}}\equiv 1-\frac{c_{r}}{p_{s^{*}}}$
(14) $p_{r_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}\ ;$ $\text{where }\alpha^{\dagger}_{r}\equiv 1-\frac{c_{r}}{p_{r^{*}}}$
@@
(15) $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}\iff\alpha^{\dagger}_{r}\leq\alpha_{s^{*},r^{*}}\iff c_{s}\leq c_{s^{*}}\ ;$ $\text{where }c_{s^{*}}\equiv\frac{c_{r}^{2}}{p_{r^{*}}}$
Finally, let’s turn our attention to the right interval of (9), $p\in\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$. Fortunately, we have already determined the relative ordering of each of the prices in the right interval, except for $p^{\dagger}$. For these conditions involving $p^{\dagger}$, we introduce the $p^{\dagger}$ relevance fees for the retailer and seller (derived in (39) and Corollary 5):
(16) $p^{\dagger}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}\ ;$ $\text{where }\alpha^{\dagger}_{r}\equiv 1-\frac{c_{r}}{p_{r^{*}}}$
@@
Note that we already introduced $\alpha^{\dagger}_{r}$ in (14). The notation arises because we will show that $\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}$ is necessary for any equilibrium where demand is fulfilled at $p^{\dagger}$. We cannot find a closed form for $\alpha^{\dagger}_{s}$ in general, though in Appendix C.4 we are still able to produce the bound $\alpha^{\dagger}_{s}\in[\max\{\alpha_{r^{*}},\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\},1-c_{s}]$. Finally, the relative ordering of $p_{s^{*}}$ and $p^{\dagger}$ is complex, and there unfortunately is no special $\hat{\alpha}$ for which $p_{s^{*}}\leq p^{\dagger}\iff\alpha\leq\hat{\alpha}$. Thus, we will leave that condition unsimplified for now.